OpenStudy (anonymous):

Rewrite with only sin x and cos x sin 3x - cos x Pwease help me

OpenStudy (anonymous):

sin 3x = sin(2x+x) = sin2x cosx +cos2x sinx now sin2x = 2sinx cosx cos2x = cos^2x -sin^2x can you do it now??

OpenStudy (anonymous):

ummm I'm a little confused still :(

OpenStudy (anonymous):

where??

OpenStudy (anonymous):

wait would the answer be 2 sin^3 x cos^4 x +1

OpenStudy (anonymous):

i dont know the answer its lengthy question but the clue is right i just gave you